# FOSP: Fine-tuning Offline Safe Policy through World Models

### v1

@@ Appendix F Additional Experimental Results
We present the training curves of FOSP during online fine-tuning in simulation experiments in Figure 8. The results show that with minimal fine-tuning, FOSP achieves better performance, outperforming SafeDreamer. Although it does not match the performance of DreamerV3, FOSP achieves nearly zero-cost of the agent throughout the entire process.
Appendix G Limitations and Future Work
Limitations
Despite achieving good results in many simulated environment experiments, the proposed framework still has some limitations. During the fine-tuning process with new trajectory data, the model struggles to update and converge efficiently with zero constraint violations. Also, we find that the model performs poorly in tasks with occluded vision. In experiments with real robots, our test results remain unstable and are sensitive to hyperparameter adjustments during training. This indicates the need for significant effort in tuning the model’s hyperparameters to achieve better performance. In transfer experiments, we find that when the positions of obstacles and target objects are random, the model has difficulty in converging with limited data despite fine-tuning. This is why we constrain the placement range of obstacles and target objects during our experiments, which also highlights the model’s high dependency on the dataset distribution.
Future Work
@@ Appendix G Limitations and Future Work
The possible future work will concentrate on three main fronts: improving safety considerations, assessing the algorithm across diverse tasks, and fortifying its robustness in real-world unseen tasks. To further decrease the violations, the safety component of our method can be improved during the offline-online stage. The SafeReach task is a preliminary experiment in the real world and it is possible to expand our method to general robot tasks such as grasping, pulling, and pushing. Moreover, the real-world task performance of our method can be further improved. To address the issue of inaccurate position determination from a single perspective and line of sight occlusion, the multi-view RL shows its efficiency in catching better features from images. What’s more, increasing the dataset capacity and uniforming the dataset distribution can probably enhance the robustness of the performance and expand it to more complex scenarios.

### v2

@@ H.1 Offline Training Results
We present the training curves of FOSP during offline pretrain in simulation experiments in Figure 13. The results show that with minimal fine-tuning, FOSP achieves better performance, outperforming SafeDreamer. Although it does not match the performance of DreamerV3, FOSP achieves nearly zero cost of the agent throughout the entire process. Note that all methods struggle to get higher rewards in SafetyPointPush because the line-of-sight obstructions necessitate online exploration.
Figure 15: Results on dataset ablation. We conducted experiments with five different dataset ratios, performing an ablation study on the proportion of safe and unsafe data. Each model was trained offline for 1 million steps and fine-tuned online for 0.8 million steps.
H.2 Dataset Ablation Results
The detailed training curves of dataset coverage ablation are shown in Figure 15. They all trained on the SafetyPointGoal2 for 1 million steps during the offline stage and 0.8 million steps for online fine-tuning. The dataset size is limited to 600 trajectories. As we can see, due to the characteristics of model-based safe RL, a high-cost, high-reward policy is learned from a purely safe dataset, while both cost and reward are low from a purely unsafe dataset. This is because the uneven distribution of the dataset leads to biases in the training of the dynamics model. Since model-based reinforcement learning algorithms rely heavily on model-generated rollouts for training, this bias can significantly impact the final decision-making. If the dataset only contains safe data, the dynamics model trained through supervised learning will mistakenly assume that the agent will always incur a cost of 0 in any state, leading the agent to ignore dangerous areas (failing to learn the cost critic). Conversely, suppose the dataset only contains unsafe data. In that case, the model will generate a large number of unsafe states that hinder the agent from completing the task, ultimately causing the agent to lose the ability to accomplish the task (overlearn the cost critic).
The dataset coverage further influences online performance. Due to offline training results, the policy pertrained on a safe dataset will have higher rewards with aggressive behaviors while the unsafe one will be more conservative on the initial stage of fine-tuning. It also affects the final performances. As a result, we adopt a compromise approach to train the policy (i.e. green curve), which achieves relatively higher rewards while maintaining zero cost.
We also illustrate the safe-random and unsafe-random mixed experiments to investigate the impact of random data in Figure 16. Compared to experiments on purely safe or purely unsafe datasets (Figure 15), adding random data can help alleviate errors in learning the world model to some extent. As shown in the training curves, experiments mixing safe data with random data resulted in lower costs during offline training than purely safe experiments. And mixing unsafe data with random data enhanced the model’s exploration, getting relatively higher rewards. Consequently, we need to add some random data to the dataset to get better performance.
Figure 16: Dataset ablation with random data. We designed a series of comparative experiments involving mixtures of safe/unsafe data and random data. Each model was trained offline for 1 million steps and fine-tuned online for 0.8 million steps.
The dataset size ablation studies are shown in Figure 17. The results align with general expectations. The policies trained on larger datasets outperform those trained on smaller datasets. However, as the amount of data increases, the improvements become less significant.
Figure 17: Dataset ablation with different sizes. We illustrate the details of dataset size ablation studies. ”Standard” represents standard size, which contains 600 trajectories. ”Larger” and ”Smaller” denote the larger one containing 900 trajectories and the smaller one containing 300 trajectories. Each model was trained offline for 1 million steps and fine-tuned online for 0.7 million steps.
H.3 Safe Generalization Results
As depicted in Figure 18, we devise more simulation experiments on safe generalization tasks. The upper one shows the performances during the offline phase. The following two figures show their generalization performance on the more challenging tasks, SafetyFadingEasy1 and SafetyFadingHard1. During the fine-tuning process, the target will gradually disappear and the agent should quickly find a path to reach the goal. The results demonstrate that FOSP can get higher rewards and lower costs even though it is deployed on tasks different from offline pretraining. Compared to SafeDreamer, it is significantly important that it can maintain near zero constraint violations during fine-tuning on new tasks.
Figure 18: More experiments on safe generalization tasks. We evaluate our method on different safe generalization tasks. The upper one illustrates the offline pretrain performances on SafetyPointGoal1. We then fine-tune the model on SafetyFadingEasy1 and SafetyFadingHard1. FOSP can consistently outperform the baseline while ensuring safe fine-tuning.
We test the specific average reward and cost for the transfer task from SafetyPointGoal1 to SafetyPointGoal2 and present the results in Figure 19. Although FOSP can outperform SafeDreamer, it may also face some costs caused by novel constraints. As shown in Figure 19, the initial cost in the new environment is slightly higher than the original but reduces quickly as fine-tuning progresses. It indicates that the safe policy expansion mechanism helps the model adapt to new safety constraints quickly. The fine-tuning process can further improve its safety performance, promoting better generalization.
Figure 19: Results on average reward and cost for transfer task. The results are evaluations on SafetyPointGoal1 to SafetyPointGoal2 task. The x-axis represents the number of fine-tuning steps. The original means the model’s performance at the end of offline pretraining in the original environment.
Table 5: Online fine-tuning and direct online results. Reward return and cost return of FOSP in offline-to-online fine-tuning (1M steps for offline pretraining) and direct online training. We report the mean value of 5 independent runs with different seeds.
Task Online fine-tuning Direct online
0 step 20k steps 50k steps 100k steps 300k steps 500k steps 500k steps 1000k steps
Reward Cost Reward Cost Reward Cost Reward Cost Reward Cost Reward Cost Reward Cost Reward Cost
PointGoal1 18.723 1.451 18.692 3.4 19.581 2.719 20.063 0.328 19.649 0.268 21.517 0.2 14.749 2.66 18.586 1.08
PointButton1 14.65 9.622 13.415 8.639 13.569 6.664 15.306 4.866 17.861 4.256 18.102 2.188 10.57 6.88 12.81 2.76
PointPush1 4.092 18.117 7.253 1.232 7.916 1.497 9.148 1.089 10.373 0.507 13.281 0.157 1.397 2.49 4.181 2.05
PointGoal2 8.1 7.556 10.31 1.486 10.157 0.427 12.446 0.18 13.034 0.214 13.556 0.234 9.471 4.779 10.717 3.089
CarGoal2 10.146 1.618 9.063 1.063 9.816 0.385 10.275 0.394 12.378 0.18 14.512 0.071 8.749 2.12 10.144 1.21
H.4 Comparisons between Online Training and Fine-tuning
To further validate the effectiveness of the fine-tuning process, we compare the results between FOSP directly trained online and FOSP with fine-tuning. We illustrate the results for different numbers of steps in the Table 5. The results denote that the online version FOSP does not perform as well as the fine-tuning method, even after training for 1M steps. Meanwhile, direct online training faces challenges in requiring a large number of training steps to converge. In contrast, our framework allows the agent to leverage offline knowledge to enhance online performance effectively with relatively short fine-tuning steps. The results also suggest that the policy can be continually improved during online fine-tuning.
H.5 More Visual Changes on Transfer Task
To evaluate the performance of our method under significant visual variations, the model pre-trained on SafetyPointGoal1 is fine-tuned in the SafetyPointBuildingGoal1 environment, with the results presented in Figure 20. The new environment has the same task as the original one but the visual inputs are different. The agent struggles to finish the task since it fails to recognize the hazardous areas and goals in the new environment. Plus, the inaccurate visual demonstrations in the offline dataset significantly impact the performance during online fine-tuning.
Based on the simulation results, we can easily know that FOSP can not deal with significant visual changes in the real world. Swapping the colors of obstacles and goals could be a straightforward setting (i.e., making obstacles red and goals green). However, some prior work has demonstrated that robots are sensitive to the color (Feng et al., 2023) and the robot is highly likely to mistakenly identify the obstacles to the goal. Additionally, it violates our original intention of using colors to distinguish obstacles and goals. Hence, the robot will fail in this setting.
Figure 20: Fine-tuning on SafetyPointBuildingGoal1. The model pretrained for 1M steps on SafetyPointGoal1 is fine-tuned on SafetyPointBuildingGoal1. While the task objectives remain the same between these two environments, the agent’s visual inputs are entirely different. In SafetyPointBuildingGoal1, the agent must avoid red-marked areas and reach the parking area (P). These hazardous areas and goals are totally different from the original ones.
H.6 Experiments on Race
We utilize a more realistic environment, Race from Ji et al. (2023), where the agent receives $64\times 64\times 3$ image inputs, as shown in Figure 21. An increase in environmental complexity better highlights the reliability of the algorithm. We employ Level 2 of the environment, which requires the agent to reach the goal position from a distant starting point while ensuring it avoids straying into the grass and prevents collisions with roadside objects. The offline dataset collected by standard procedures is a mixture of safe, unsafe and random data.
As illustrated in Figure 22, FOSP has superior performance and lower constraint violations during the offline training phase. In the online fine-tuning stage, it further optimizes safety performance while ensuring a stable improvement in task rewards. Conversely, SafeDreamer exhibits higher constraint violations during the offline phase, leading to unsafe behaviors at the beginning of the online fine-tuning. Additionally, its safety design limitations make it challenging to achieve nearly zero constraint violations.
Figure 21: Race environment. The left subfigure shows the panoramic picture of the environment. The right subfigure is the first perspective of the agent.
Figure 22: Race experiments results. We compare FOSP with SafeDreamer on Race2. Each model was trained offline for 1 million steps and online for 0.5 million steps. FOSP has comparable rewards but lower costs than SafeDreamer.
Appendix I Limitation and Future Work
I.1 Limitations
Safe constraints
Despite achieving good results in many simulated environment experiments, the proposed framework still has some safety concerns. First, it inevitably requires a balanced sampled and sufficiently large dataset for offline policy pertraining, which can be undesirable for safety-critical applications. An uneven dataset, such as one lacking unsafe data, or an insufficiently large dataset, can lead to deviations in model learning, thereby compromising its safety performance and generalization. Second, it may exceed the feasible region from time to time if the cost distribution has a long tail since CMDP only requires the policy to satisfy the expectation cost constraint. It might be considered that incorporating risk-constrained methods (Chow et al., 2018) can solve problems with long tail cost distributions.
Real-world experiments
In experiments with real robots, we find that the model performs poorly in tasks with occluded vision. Our test results remain unstable and are sensitive to learning rate adjustments. This indicates the need for significant effort in tuning the model’s hyperparameters to achieve better performance. In addition, SafeReach task is quite preliminary. It is difficult to apply FOSP to more realistic scenarios like real-time control or dynamic obstacles as it lacks certain predictive capabilities. The model also fails to handle novel visual observations beyond scene re-configurations during fine-tuning because it struggles to interpret the meaning of different objects in new environments.
I.2 Future Works
More complex real-world tasks
To further decrease the violations, the safety component of our method can be improved during the offline-online stage. The SafeReach task is a preliminary experiment in the real world and it is possible to expand our method to general robot tasks such as grasping, pulling, and pushing. For the tasks with novel visual observations, leveraging semantic information into the inputs is a promising solution, which can help the agent attach meaningful interpretations to raw sensory inputs.
Improvement in real-world scenarios
@@ I.2 Future Works
The real-world task performance of our method can be further improved. To address the issue of inaccurate position determination from a single perspective and line of sight occlusion, the multi-view RL shows its efficiency in catching better features from images. Increasing the dataset capacity and uniforming the dataset distribution can probably enhance the robustness of the performance and expand it to more complex scenarios.
Safe sim2real
FOSP also shows potential in the field of safe sim-to-real transfer. Due to the limitations of simulators in fully replicating real-world environments, robots may encounter safety challenges when deploying algorithms in the real world. Thus, it is important to fine-tune it in the real world with safety considerations. Meanwhile, the use of world models will help speed up the fine-tuning process.